two vector have a magnitude F and 2root F .they are inclined to each other at an angle of 135 degree .find their resultant in magnitude and direction .how to I solve it
1
Expert's answer
2020-06-02T17:52:31-0400
"R=\sqrt{F^2+(\sqrt{2}F)^2+2F\sqrt{2}F\cos{135}}"

"R=\sqrt{1+2-2\sqrt{2}\cos{45}}F=F"

For direction you should have to use this formula


"\tan{\alpha}=\frac{F\sin{135}}{F\cos{135}+F}=\frac{\sin{135}}{\cos{135}+\sqrt{2}}"

"\alpha=45\degree"